### v3

# Point2RBox-v3: Self-Bootstrapping from Point Annotations via Integrated Pseudo-Label Refinement and Utilization

Abstract
Driven by the growing need for Oriented Object Detection (OOD), learning from point annotations under a weakly-supervised framework has emerged as a promising alternative to costly and laborious manual labeling. In this paper, we discuss two deficiencies in existing point-supervised methods: inefficient utilization and poor quality of pseudo labels. Therefore, we present Point2RBox-v3. At the core are two principles: 1) Progressive Label Assignment (PLA). It dynamically estimates instance sizes in a coarse yet intelligent manner at different stages of the training process, enabling the use of label assignment methods. 2) Prior-Guided Dynamic Mask Loss (PGDM-Loss). It is an enhancement of the Voronoi Watershed Loss from Point2RBox-v2, which overcomes the shortcomings of Watershed in its poor performance in sparse scenes and SAM’s poor performance in dense scenes. To our knowledge, Point2RBox-v3 is the first model to employ dynamic pseudo labels for label assignment, and it creatively complements the advantages of SAM model with the watershed algorithm, which achieves excellent performance in both sparse and dense scenes. Our solution gives competitive performance, especially in scenarios with large variations in object size or sparse object occurrences: 66.09%/56.86%/41.28%/46.40%/19.60%/45.96% on DOTA-v1.0/DOTA-v1.5/DOTA-v2.0/DIOR/STAR/RSAR.
[…]
1 INTRODUCTION
Oriented object detection (OOD) has attracted growing attention due to the increasing demand for object direction estimation in diverse fields, including autonomous driving [feng2021deep], aerial images [fu2020rotation, liu2017high, xia2018dota, yang2022arbitrary, yang2018automatic], scene text [liao2018rotation, liu2018fots, zhou2017east], retail [goldman2019precise, pan2020dynamic], and industrial inspection [liu2020data, wu2022pcbnet].
Training OOD models requires annotations as supervisory signals. Traditional rotated bounding boxes (RBoxes) provide accurate supervision but are costly: annotating each RBox is 36.5% more expensive than a horizontal box (HBox) and 104.8% more than a point [yu2024point2rbox], highlighting the potential of weakly-supervised OOD. HBox-based approaches such as H2RBox [yang2023h2rbox] and H2RBox-v2 [yu2023h2rboxv2, yu2025wholly] have shown strong results, while point-based methods are rapidly advancing.
Current point-supervised methods fall into four categories: (1) Pseudo generation using multiple instance learning and class probability maps [luo2024pointobb, pointobbv2]; (2) Knowledge combination with one-shot learning [yu2024point2rbox]; (3) Point-prompt OOD leveraging SAM’s zero-shot capability [cao2023p2rbox, 10649581, liu2025pointsam, lu2025semantic, kirillov2023segany]; (4) Spatial layout for pseudo label generation [yu2025point2rbox]. Despite these advances, pseudo label quality remains a bottleneck, which motivates our approach.
Motivation. Our motivation mainly stems from the lack of quality and utilization efficiency of pseudo labels in all current end-to-end methods. End-to-end point supervision method requires label assignment for the Feature Pyramid Network (FPN) [lin2017feature], which plays a crucial role in terms of detecting scale information. Most existing methods simply distribute them to the same layer, such as Point2RBox-v2 [yu2025point2rbox]. This wastes the scale information contained in the pseudo labels generated by the model. Point2RBox-v2 benefits from the “Voronoi Watershed Loss” for generating masks as pseudo labels. However, the watershed algorithm is less effective in sparse scenes. As SAM can also provide masks, we investigated it as an alternative but found that SAM fails in dense scenes. Can we enhance the detection capability of the model by focusing on these two aspects? In this paper, we have an in-depth discussion about this issue to explore how to fully uncover the potential of pseudo labels.
Highlights. 1) Point2RBox-v3 is proposed for point-supervised OOD, advancing the state of the art (SOTA) as displayed in Figure 1 and Tables 1-2. 2) We improve the quality and utilization efficiency of pseudo labels, compared to Point2RBox-v2, without incurring a significant increase in training costs, which reveals more possibilities for the application and quality of pseudo labels in weakly-supervised object detection for future researchers.
Contributions. 1) To our knowledge, Point2RBox-v3 is the first end-to-end point supervision model to explore how to assign pseudo labels to multiple FPN levels, and it complements the advantages of the SAM model with the watershed algorithm, which achieves excellent performance in both sparse and dense scenes. 2) We extend our method to partial weakly-supervised tasks [liu2025partial] beyond point supervision, demonstrating its adaptability and scalability. 3) The training pipeline and detailed implementation are elucidated. The source code will be made publicly available.
Figure 1: (Left) Visual comparisons with the state-of-the-art method Point2RBox-v2. The boxes detected by our method wrap the objects more tightly, with fewer missed detections. (Right) Radar plot comparing the performance of our method with 10 other state-of-the-art methods across 6 benchmark datasets. Methods with ‘(e)’ in the legend indicate end-to-end version, while those without ‘(e)’ represent two-stage version.
2 Related Work
Figure 2: The training pipeline of Point2RBox-v3. Progressive Label Assign utilizes scale info from pseudo label to dynamically assign gt point (see Figure 3). Prior-Guided Dynamic Mask provides enhanced mask supervision information (see Figure 4). $\mathcal{L}_{others}$ are the loss functions inherited from Point2RBox-v2 (see Appendix 6.2).
Point-supervised Oriented Detection. Due to the low-cost and high-quality requirements in the field of object detection, point-supervised oriented object detection has become one of the research focuses. Some models apply the powerful zero-shot segmentation performance of the Segment Anything Model (SAM) [kirillov2023segany] in their pipelines, such as P2RBox [cao2023p2rbox], PMHO [zhang2024pmho], PointSAM [liu2024pointsam] and PMS-SAM-RSD [lu2025semantic]. Apart from these models, there are also some other models whose strategies involve creating pseudo labels using traditional machine learning or other methods. For example, the PointOBB series [luo2024pointobb, ren2024pointobbv2, zhang2025pointobbv3] employs techniques such as multi-instance learning and class probability graphs to generate pseudo RBoxes. Differently, Point2RBox [yu2024point2rbox] acquires knowledge from single sample examples through a knowledge combination method, and Point2RBox-v2 [yu2025point2rbox] applies traditional machine learning algorithms such as watershed and edge detection. The technology based on SAM can achieve simpler applications by leveraging powerful base models, while other methods offer a solution that does not require reliance on other components. However, existing approaches have not yet explored their combination to enhance model performance.
Label Assignment. Point-supervised oriented object detection lacks scale information, and therefore cannot directly apply label assignment. PointSAM [liu2025pointsam] adopts a ViT [dosovitskiy2020image] backbone. Point2RBox [yu2024point2rbox], Point2RBox-v2 [yu2025point2rbox] adopt only a single FPN feature level. PointOBB [zhang2025pointobb] selects grid points within the central region around each ground-truth point across all feature levels as positive samples. PointOBB-v3 [zhang2025pointobbv3] applies a gating mechanism to each FPN feature layer during training to produce corresponding gating scores, which are then used to automatically aggregate the multiple FPN output layers into a fused feature map for feature extraction. Unfortunately, these methods discard the classical multi-level label assignment in FPNs. We show that this omission largely explains the gap between point and full supervision, and that leveraging coarse scale cues from pseudo labels can effectively narrow this gap.
3 Method
Figure 3: The process of Progressive Label Assignment (PLA). Points of different colors represent those assigned to different feature pyramid levels P2, P3, P4, P5, P6 for label assignment. As training progresses, the label assignment strategy evolves. It begins with using fixed Watershed regions in the early stages and transitions to leveraging dynamic, network-generated dimensions in the middle-to-late phases. This evolution guides ground truth points to be assigned to more suitable FPN levels over time.
3.1 Overview and Preliminary
In brief, the definition of the Point-supervised Object Detection task is as follows: during the training phase, the model inputs consist of an image $I$ and the point annotations $P=\{(x_{i},y_{i})\}$ for all instances within the image (these points are typically, though not strictly, defined as the center points of instances). The training objective is to output the Rotated Bounding Box (RBox) representation $(x,y),(w,h),\theta]$ and the instance category $[cls]$ for each instance, where $(x,y),(w,h),\theta]$ denotes the precise circumscribed rectangle of the instance. During the inference phase, the input is solely the image $I$, and the model outputs the RBoxes $(x,y),(w,h),\theta]$ of all potential instances in the image, along with their corresponding categories $[cls]$ and confidence scores. The core difficulty of this task lies in overcoming the extreme deficiency of scale and orientation information inherent in point annotations. Specifically, the model is required to “reconstruct” precise geometric information from scratch. This requires resolving mutual boundary interference in densely packed scenes and addressing segmentation difficulties in sparse scenes caused by the scarcity of spatial constraints, while overcoming the challenge of diversity in scales.
The architectural overview of Point2RBox-v3 is presented in Figure 2. The simple evolutionary context of Point2RBox series is as follows: Point2RBox [yu2024point2rbox] established the angle prediction module via symmetry learning, while Point2RBox-v2 [yu2025point2rbox] introduced the scale prediction module based on spatial layout learning. The detailed explanation is in Appendix 6.1. While retaining the foundational components of this strong baseline (i.e., ResNet50 [he2016deep] backbone, FPN [lin2017feature] head, PSC [yu2023psc] angle coder, and losses $\mathcal{L}_{others}$ as shown in the Figure 2), Point2RBox-v3 focuses on two core innovations: Progressive Label Assignment (PLA) and Prior-Guided Dynamic Mask Loss (PGDM-Loss). These modules not only reinforce scale learning but, crucially, PLA enables the utilization of dynamic scale information for multi-level label assignment within the FPN under a weakly-supervised framework.
PLA iteratively generates and refines online pseudo bounding boxes in a dynamic manner for label assignment. It re-establishes the effectiveness of FPN in point-supervised learning scenarios. PGDM-Loss enhances the supervisory quality in the object scale learning task by selecting the mask-generating method between SAM and watershed. Incorporated with other components inherited from Point2RBox-v2, our method achieves a consistent improvement (59.6% vs 51.0%) over the previous SOTA end-to-end performance. In subsequent subsections, PLA and PGDM-Loss are detailed.
3.2 Progressive Label Assignment
For both anchor-free [tian2019fcos] and anchor-based [lin2017focal] object detection algorithms, the scale information of objects plays an indispensable role in label assignment. Points lack scale, rendering label assignment infeasible. The scarcity of scale information persists and is even magnified in the loss constraints for scale learning. In Point2RBox-v2, the authors employ various approaches for scale learning, including explicit methods such as copy-paste loss, Voronoi Watershed loss, and edge loss, as well as implicit techniques like layout loss. Among these, watershed segmentation algorithm plays the most pivotal role in the scale learning task of the entire framework. Inspired by this, we introduce the pseudo labels originally used for scale constraints in the loss function into the label assignment module to provide approximate scale information.
We re-adopted the standard Feature Pyramid Network (FPN) architecture and the label assignment strategy. In the initial phase of model training, scale cues are obtained from watershed-generated pseudo labels. The computation of pseudo labels can be represented by the formula:
$V=\text{Voronoi}(X),$ (1)
$S=\text{Watershed}(I,X,V),$ (2)
$PL=\text{minAreaRect}(S),$ (3)
where Voronoi [aurenhammer1991voronoi] is a partitioning of a space based on a set of points; $X$ are the annotated points within a training image; $I$ is the input image; $V$ are the output ridges; $S$ are the output basin regions (pixel coordinates) corresponding to each annotated instance; minAreaRect is used to calculate the minimum enclosing rotated rectangle of a point set; $PL$ is for Pseudo Label.
Static segmentation algorithms such as watershed produce immutable segmentation regions. When such algorithms yield suboptimal segmentation for specific samples, the resulting defects persist throughout the entire training cycle without correction. Consequently, during the mid-to-late training phases, we utilize the forward network-predicted boxes associated with each GT point to supply requisite dimension information for label assignment. For each feature level in the FPN, it selects the predicted boxes from the anchor point closest to the target point as the candidate boxes. The final pseudo label is then chosen from these candidate boxes based on their corresponding scores. The computation of pseudo labels can be represented by the formula:
$PL_{g}=\underset{b\in C_{g}}{\arg\max}{score}(b)$ (4)
where $PL_{g}$ is the Pseudo Label for ground truth point $g$; $C_{g}$ refers to the candidate prediction boxes, selected by picking the prediction box associated with the anchor point closest to the ground-truth point $g$ at each FPN level; function $score(b)$ is the raw classification confidence directly output by the detection network. This approach can partially mitigate label assignment issues caused by solely relying on fixed segmentation regions generated by the watershed algorithm.
The improvement of the above design is verified in the Table 5 of the ablation study. Algorithm 1 in Appendix 6.3 outlines the workings of the proposed method. Figure 3 illustrates the limitations of using fixed segmentation regions, such as those generated by watershed, for label assignment scale information. It also shows how, in the later stages of training, the dynamic forward predictions of the network progressively improve label assignment accuracy.
3.3 Prior-Guided Dynamic Mask Loss
Figure 4: Comparison between watershed and SAM masks on DOTA-v1.0. The red patches with yellow edges represent the masks generated by the model. The processing result in the top-right corner shows significant over-segmentation by SAM, which causes the masks to visually merge into a large, incorrect region.
Point2RBox-v2 [yu2025point2rbox] leverages a Watershed Loss to generate pseudo labels from spatial layouts. While effective in dense scenes, this approach struggles in sparse scenarios where spatial cues are minimal, often leading to over- or under-segmentation. A detailed analysis of this limitation is provided in Appendix 6.4. Conversely, the Segment Anything Model (SAM) [kirillov2023segany] offers greater robustness in these challenging scenarios, but its high computational cost and weaker performance in dense scenes [cai2024crowd] preclude its direct application.
To harness the strengths of both methods, we propose the PGDM-Loss ($\mathcal{L}_{PGDM}$), a hybrid loss that dynamically routes images for mask generation. Images with a sparse instance count (total instances $\leq N_{thr}$) are directed to a SAM branch, while denser scenes are processed by the original -Watershed branch (see Figure 4). This strategy enhances segmentation accuracy for difficult cases without a significant computational overhead, preserving the efficiency of the Voronoi Watershed method in dense scenes. We choose the lightweight MobileSAM [mobile_sam] as our SAM model. On DOTA-v1.0, its AP50 is the same as the result obtained using the basic SAM model, but its training time is 10 hours less than that of the basic one.
We integrate SAM not as a simple mask proposal tool, but as a source of weak supervision. Therefore, the SAM model does not participate in the inference process at all, ensuring its fast inference speed. For an instance $j$ of class $c_{j}$ routed to the SAM branch, SAM generates a set of candidate masks, $M_{j}=\{m_{1},m_{2},\dots,m_{k}\}$. We select the optimal mask $m_{j}^{*}$ via a prior-guided filtering mechanism governed by a scoring function:
$m_{j}^{*}=\underset{m_{i}\in M_{j}}{\arg\max}\sum_{k}w_{k,c_{j}}\cdot\phi_{k}(m_{i}),$ (5)
where $\phi_{k}(m_{i})$ represents metrics calculated from the masks, and $w_{k,c_{j}}$ is a class-specific weight based on simple prior knowledge (e.g., expected shape). A full description of the five metrics (center alignment, color consistency, rectangularity, circularity, and aspect ratio reliability) is available in Appendix 6.5.
After obtaining masks from both branches, we compute the losses uniformly, following Point2RBox-v2 [yu2025point2rbox]. First, the mask $S$ is rotated to align with the current prediction to obtain the regression targets $w_{t}$ and $h_{t}$:
[…]
The width-height regression loss for a single instance, which we denote as $\mathcal{L}_{mask}$, is then computed using the Gaussian Wasserstein Distance loss ($\mathcal{L}_{GWD}$) [yang2021rethinking]:
$\mathcal{L}_{mask}=\mathcal{L}_{GWD}\left(\begin{pmatrix}w/2&0\\ 0&h/2\end{pmatrix}^{2},\begin{pmatrix}w_{t}/2&0\\ 0&h_{t}/2\end{pmatrix}^{2}\right),$ (7)
where $w$ and $h$ is the output of the detection head. The total loss, $\mathcal{L}_{PGDM}$, is the mean of these individual losses over all $N$ instances in an image. We denote the loss for instance $j$ as $\mathcal{L}_{j}$ and formulate the total loss as:
$\mathcal{L}_{PGDM}=\frac{1}{N}\sum_{j\in Image}\mathcal{L}_{j},$ (8)
where each $\mathcal{L}_{j}$ is computed for the corresponding instance using its optimal mask according to the formula for $\mathcal{L}_{mask}$ in Eq. 7.
4 Experiments
4.1 Experimental Details
Implementation Details. Experiments are carried out using PyTorch 1.13.1 [paszke2019pytorch] and the rotation detection toolkit MMRotate 1.0.0 [zhou2022mmrotate]. All experiments follow the same hyper-parameters. We adopt Average Precision (AP) as the primary evaluation metric. All models are based on the ResNet50 [he2016deep] backbone and are trained using AdamW [loshchilov2018decoupled].
  1. 1.
    Learning rate: Initialized at $5\times 10^{-5}$, with a warm-up for 500 iterations, and divided by 10 at each decay step.
  2. 2.
    Epochs: 12 for all datasets.
  3. 3.
    Augmentation: Random flip for all datasets.
  4. 4.
    Image size: Images from DOTA/STAR are split into 1,024 × 1,024 patches with an overlap of 200 pixels; images from DIOR are scaled to 800 × 800; images from RSAR are scaled to 1024 × 1024.
  5. 5.
    Multi-scale: All experiments are evaluated without multi-scale techniques [zhou2022mmrotate].
Datasets. The experiments utilize six remote sensing datasets and one retail scene dataset, covering those used by the main counterparts:
  • DOTA [xia2018dota]. DOTA-v1.0 contains 2,806 aerial images with 15 categories. DOTA-v1.5/2.0 are its extended versions with more small objects and new categories.
  • DIOR [li2020object]. It is an aerial image dataset re-annotated with RBoxes based on its original HBox version [cheng2022anchor], featuring high variation in object scale and high intra-class diversity.
  • STAR [li2024star]. An extensive dataset for scene graph generation, covering over 210,000 objects with diverse spatial resolutions and 48 fine-grained categories.
  • RSAR [zhang2025rsar]. A remote sensing dataset based on Synthetic Aperture Radar (SAR) imagery, containing 6 categories.
4.2 Main Results on DOTA-v1.0
[…]
RepPoints [yang2019reppoints] 86.7 81.1 41.6 62.0 76.2 56.3 75.7 90.7 80.8 85.3 63.3 66.6 59.1 67.6 33.7 68.45
RetinaNet [lin2017focal] 88.2 77.0 45.0 69.4 71.5 59.0 74.5 90.8 84.9 79.3 57.3 64.7 62.7 66.5 39.6 68.69
GWD [yang2021rethinking] 89.3 75.4 47.8 61.9 79.5 73.8 86.1 90.9 84.5 79.4 55.9 59.7 63.2 71.0 45.4 71.66
FCOS [tian2019fcos] 89.1 76.9 50.1 63.2 79.8 79.8 87.1 90.4 80.8 84.6 59.7 66.3 65.8 71.3 41.7 72.44
S2A-Net [han2022align] 89.2 83.0 52.5 74.6 78.8 79.2 87.5 90.9 84.9 84.8 61.9 68.0 70.7 71.4 59.8 75.81
[…]
Sun et al. [sun2021oriented] $\times$ 51.5 38.7 16.1 36.8 29.8 19.2 23.4 83.9 50.6 80.0 18.9 50.2 25.6 28.7 25.5 38.60
BoxInst-RBox [tian2021boxinst]2 $\times$ 68.4 40.8 33.1 32.3 46.9 55.4 56.6 79.5 66.8 82.1 41.2 52.8 52.8 65.0 30.0 53.59
H2RBox [yang2023h2rbox] 88.5 73.5 40.8 56.9 77.5 65.4 77.9 90.9 83.2 85.3 55.3 62.9 52.4 63.6 43.3 67.82
EIE-Det [wang2024explicit] 87.7 70.2 41.5 60.5 80.7 76.3 86.3 90.9 82.6 84.7 53.1 64.5 58.1 70.4 43.8 70.10
H2RBox-v2 [yu2023h2rboxv2] 89.0 74.4 50.0 60.5 79.8 75.3 86.9 90.9 85.1 85.0 59.2 63.2 65.2 70.5 49.7 72.31
[…]
Point2Mask-RBox [li2023point2mask]2 $\times$ 4.0 23.1 3.8 1.3 15.1 1.0 3.3 19.0 1.0 29.1 0.0 9.5 7.4 21.1 7.1 9.72
P2BNet+H2RBox [yang2023h2rbox] $\times$ 24.7 35.9 7.1 27.9 3.3 12.1 17.5 17.5 0.8 34.0 6.3 49.6 11.6 27.2 18.8 19.63
P2BNet+H2RBox-v2 [yu2023h2rboxv2] $\times$ 11.0 44.8 14.9 15.4 36.8 16.7 27.8 12.1 1.8 31.2 3.4 50.6 12.6 36.7 12.5 21.87
P2RBox [yu2024point2rbox] $\times$ 87.8 65.7 15.0 60.7 73.0 71.7 78.9 81.5 44.5 81.2 41.2 39.3 45.5 57.5 41.2 59.04
PointOBB [luo2024pointobb] $\times$ 26.1 65.7 9.1 59.4 65.8 34.9 29.8 0.5 2.3 16.7 0.6 49.0 21.8 41.0 36.7 30.08
Point2RBox [yu2024point2rbox] 62.9 64.3 14.4 35.0 28.2 38.9 33.3 25.2 2.2 44.5 3.4 48.1 25.9 45.0 22.6 34.07
Point2RBox+SK [yu2024point2rbox] 53.3 63.9 3.7 50.9 40.0 39.2 45.7 76.7 10.5 56.1 5.4 49.5 24.2 51.2 33.8 40.27
Point2RBox+SK [yu2024point2rbox] $\times$ 66.4 59.5 5.2 52.6 54.1 53.9 57.3 90.8 3.2 57.8 6.1 47.4 22.9 55.7 40.5 44.90
PointOBB-v2 [ren2024pointobbv2] $\times$ 64.5 27.8 1.9 36.2 58.8 47.2 53.4 90.5 62.2 45.3 12.1 41.7 8.1 43.7 32.0 41.68
PMS-SAM-RSD [lu2025semantic] 69.0 39.5 6.7 44.8 64.7 71.9 79.6 79.8 2.7 60.0 12.1 32.6 39.6 44.8 42.5 46.00
PointOBB-v3 [zhang2025pointobbv3] 30.9 39.4 13.5 22.7 61.2 7.0 43.1 62.4 59.8 47.3 2.7 45.1 16.8 55.2 11.4 41.29
PointOBB-v3 [zhang2025pointobbv3] $\times$ 52.9 54.4 21.3 52.7 65.6 44.9 67.8 87.2 26.7 73.4 32.6 53.3 39.0 56.4 10.2 49.24
Point2RBox-v2 [yu2025point2rbox] 78.4 52.7 8.3 40.9 71.0 60.5 74.7 88.7 65.5 72.1 24.4 26.1 30.1 50.7 21.0 51.00
Point2RBox-v2 [yu2025point2rbox] $\times$ 88.0 72.6 8.0 46.2 79.6 76.3 86.9 89.1 79.7 82.9 26.2 45.3 45.8 66.3 46.3 62.61
Point2RBox-v2+PLA (ours) 82.3 47.8 23.2 36.0 77.9 75.5 86.6 85.3 72.0 76.5 28.0 33.4 39.4 52.7 31.7 56.55
Point2RBox-v2+PLA (ours) $\times$ 88.3 71.5 26.3 45.9 78.8 76.9 87.5 86.9 74.6 83.1 47.7 45.2 49.0 60.7 46.7 64.61
Point2RBox-v2+PLA+CS3 (ours) 84.6 51.5 25.4 37.2 78.7 75.3 86.5 86.4 75.8 74.9 31.3 33.6 40.2 54.5 34.2 58.00
Point2RBox-v3 (ours) 86.5 53.4 35.7 37.5 78.8 75.3 86.3 86.6 66.0 80.2 29.7 49.9 36.7 59.1 32.5 59.61
Point2RBox-v3 (ours) $\times$ 89.0 72.5 41.6 45.1 79.4 76.4 87.7 85.1 75.7 84.8 44.4 55.4 45.7 59.8 48.8 66.09
Comparison tracks: ✓= End-to-end training and testing; $\times$ = Generating pseudo labels to train the FCOS detector (two-stage training).
P2RBox/PMS-SAM-RSD/Point2RBox-v3: Pre-trained SAM model; Point2RBox+SK: One-shot sketches for each class.
1PL: Plane, BD: Baseball diamond, BR: Bridge, GTF: Ground track field, SV: Small vehicle, LV: Large vehicle, SH: Ship, TC: Tennis court,
  BC: Basketball court, ST: Storage tank, SBF: Soccer-ball field, RA: Roundabout, HA: Harbor, SP: Swimming pool, HC: Helicopter.
2RBox: The minimum rectangle operation is performed on the output Mask to obtain the RBox.
3CS: CS denotes the abbreviation for Class-Specific Watershed. The Appendix 6.6 contains detailed information on Class-Specific Watershed.
The best score is in bold and the second-best is in underline.
Table 1 compares Point2RBox-v3 with SOTA methods. These methods can be categorized into two tracks:
  1. 1.
    End-to-end training. These methods apply the trained weakly-supervised detector directly to the test set. Our method, demonstrates an improvement of $8.61\%$ ($59.61\%$ vs. $51.00\%$). Notably, our method also outperforms P2RBox ($59.04\%$), which also uses a pre-trained SAM.
  2. 2.
    Two-stage training. In this track, RBox pseudo labels are generated on the train/val sets, which are then used to train a standard FCOS detector. In this mode, Point2RBox-v3 achieves an accuracy of $66.09\%$, considerably surpassing the PointOBB series. Compared to its predecessor, Point2RBox-v2 ($62.61\%$), our method shows an improvement of $3.48\%$. It also outperforms the SAM-powered P2RBox ($59.04\%$) by $7.05\%$.
Class-wise Analysis. Our method demonstrates robust performance on high-density categories (SH, SV, LV, PL, ST, TC), matching the strong Point2RBox-v2 baseline. More significantly, it achieves notable gains on large-sized, low-density categories—Soccer-ball field (SBF), Bridge (BR), and Roundabout (RA)—where traditional point-supervised methods often fail to infer accurate rotated boxes from sparse points. This leads to state-of-the-art results on these challenging categories, exemplified by a dramatic improvement on BR (41.6% vs. 8.0%), strong performance on RA (55.4%), and a solid result on SBF (44.4%).
4.3 Results on More Datasets
Table 2: AP50 comparisons on the DOTA-v1.0/1.5/2.0, DIOR, STAR, and RSAR datasets.
Methods * DOTA-v1.0 DOTA-v1.5 DOTA-v2.0 DIOR STAR RSAR
[…]
RetinaNet [lin2017focal] 68.69 60.57 47.00 54.96 21.80 57.67
GWD [yang2021rethinking] 71.66 63.27 48.87 57.60 25.30 57.80
FCOS [tian2019fcos] 72.44 64.53 51.77 59.83 28.10 66.66
S2A-Net [han2022align] 75.81 66.53 52.39 61.41 27.30 66.47
[…]
Sun et al. [sun2021oriented] $\times$ 38.60 - - - - -
H2RBox [yang2023h2rbox] 70.05 61.70 48.68 57.80 17.20 49.92
H2RBox-v2 [yu2023h2rboxv2] 72.31 64.76 50.33 57.64 27.30 65.16
AFWS [lu2024afws] 72.55 65.92 51.73 59.07 - -
[…]
P2RBox [yu2024point2rbox] $\times$ 59.04 - - - - -
PMHO [zhang2024pmho] $\times$ - - - 46.20 - -
PointOBB [luo2024pointobb] $\times$ 30.08 10.66 5.53 37.31 9.19 13.80
Point2RBox+SK [yu2024point2rbox] 40.27 30.51 23.43 27.34 7.86 27.81
PointOBB-v2 [ren2024pointobbv2] $\times$ 41.68 30.59 20.64 39.56 9.00 18.99
PMS-SAM-RSD [lu2025semantic] 46.00 - - - - -
PointOBB-v3 [zhang2025pointobbv3] 41.20 31.25 22.82 37.60 11.31 15.84
PointOBB-v3 [zhang2025pointobbv3] $\times$ 49.24 33.79 23.52 40.18 12.85 22.60
Point2RBox-v2 [yu2025point2rbox] 51.00 39.45 27.11 34.70 7.80 28.60
Point2RBox-v2 [yu2025point2rbox] $\times$ 62.61 54.06 38.79 44.45 14.20 30.90
Point2RBox-v3 (ours) 59.61 48.34 34.47 41.50 14.60 40.80
Point2RBox-v3 (ours) $\times$ 66.09 56.86 41.28 46.40 19.60 45.96
* Comparison tracks: $\checkmark$ = End-to-end training and testing; $\times$ = Generating pseudo labels to train the FCOS detector (two-stage training). P2RBox/PMS-SAM-RSD/PMHO/Point2RBox-v3: Pre-trained SAM model; Point2RBox+SK: One-shot sketches for each class. The best score is in bold and the second-best is in underline.
As shown in Table 2, Point2RBox-v3 also performs excellently on other datasets. On the more challenging DOTA-v1.5/2.0 datasets, Point2RBox-v3 shows a similar leading trend. In the end-to-end track, its AP50 is $17.09\%$ and $11.65\%$ higher than PointOBB-v3, respectively, and it consistently outperforms its predecessor, Point2RBox-v2 ($56.86\%$ / $41.28\%$ vs. $54.06\%$ / $38.79\%$). On the DIOR dataset, where object distribution is relatively sparse, our method achieves an AP50 of $46.40\%$. This result is slightly higher than PMHO ($46.20\%$), whose five-stage “point - proposal bag - mask - horizontal box - rotated box” pipeline makes it expensive to train and susceptible to error accumulation, while evidently higher than other methods. On the fine-grained dataset STAR, our method achieves a competitive AP50 of $19.60\%$. Furthermore, on the SAR image dataset RSAR, Point2RBox-v3 obtains an AP50 of $45.96\%$, demonstrating its effectiveness across different imaging modalities.
4.4 Ablation Studies
Table 3: Ablation with addition of modules.
Modules DOTA
PLA PGDM E2E FCOS
51.0 62.6
56.6 64.6
54.2 63.9
59.6 66.1
Table 4: Ablation with switch epoch of PLA.
switch epoch E2E FCOS
0 56.6 64.4
3 59.5 66.4
6 59.6 66.1
9 56.2 65.5
12 56.3 65.3
Table 5: Ablation on sparse scene threshold in PGDM-Loss.
Threshold E2E FCOS Time
0 56.6 64.6 13.6h
4 59.6 66.1 19.5h
8 58.9 66.1 23.5h
$\infty$ 57.2 66.1 79.0h
Tables 5- 7 display the ablation studies on DOTA-v1.0. “E2E” denotes end-to-end training; “FCOS” denotes two-stage training (i.e. generating pseudo labels to train FCOS). The final values adopted are highlighted in gray.
Incremental addition of modules. Table 5 demonstrates the impact of different modules on the final result. Adding only the PLA module increases the E2E and two-stage metrics by 5.6% and 2.0%, reaching 56.6% and 64.6%, respectively. Adding only the PGDM-Loss module enhances these metrics by 3.2% and 1.3%, achieving 54.2% and 63.9%, respectively. The improvements from the two modules exhibit a high degree of orthogonality. When integrated simultaneously, they yield a combined gain of 8.6% and 3.5% on the E2E and two-stage metrics, respectively.
Switch epoch of PLA. Table 5 studies the hyperparameter switch epoch of PLA. Our ablation study confirms that the phased assignment strategy is crucial. The extreme methods—using only network predictions (switch epoch = 0) or only watershed boxes (switch epoch = 12) for the entire training—performed substantially worse than a mid-training transition. The optimal switch epoch was found to be 3 or 6; we selected 6 for subsequent experiments.
Hyperparameters of PGDM-Loss. We investigated the hyperparameter $N_{thr}$, a threshold on the number of target instances used to define sparse scenes. A scene is classified as sparse if its instance count is no more than $N_{thr}$. As shown in Table 5, the model achieves optimal E2E performance of 59.6% at $N_{thr}=4$ and their two-stage training performance on the FCOS platform is very close. This indicates that the model is robust to the choice of $N_{thr}$ within a reasonable range, obviating the need for meticulous tuning. Notably, setting $N_{thr}$ to infinity, which directs all instances to the SAM branch, quadruples the training time and significantly degrades precision. This result, consistent with our findings in Figure 4, underscores the importance of our hybrid approach that applies SAM and watershed strategies to different scenes based on sparsity.
Necessity of Prior Knowledge in PGDM-Loss. It is crucial to note that the reliance on prior knowledge stems from the fact that SAM is trained on general-purpose datasets. We find that while SAM can effectively segment instances based on edge features in remote sensing scenes, its native confidence scores often fail to accurately reflect the quality of the generated masks in this specific domain. For example, in Figure 2, the ‘Prior-Guided Selector’ part shows that SAM will not tend to give the most correct answer a higher score. To validate this, we conducted an ablation study on DOTA-v1.0 as shown in Table 7. The results indicate that relying solely on SAM’s native confidence scores leads to a performance degradation, with AP50 dropping by 1.75 and 2.5 points for the end-to-end and two-stage frameworks, respectively. This empirical evidence strongly demonstrates the superiority of our proposed method.
Table 6: Ablation with Prior Knowledge in PGDM-Loss
Modules E2E FCOS
NoPrior-Loss 57.86 63.59
PGDM-Loss 59.61 66.09
Table 7: AP50 comparison on DOTA-v1.0/v1.5
Method DOTA-v1.0 DOTA-v1.5
10% 20% 30% full 10% 20% 30% full
PWOOD 42.35 45.01 49.12 55.87 35.33 41.54 43.02 46.83
PGDM 42.81 47.70 51.02 57.96 37.29 42.02 44.06 49.08
PLA 45.18 50.33 55.67 62.41 40.36 45.24 48.13 53.23
PLA + PGDM 50.67 52.22 58.02 64.57 43.25 47.05 50.51 56.18
We integrated this method into the “Partially Weakly Supervised Oriented Object Detection”, i.e. PWOOD framework [liu2025partial], to demonstrate the universality and practicality of our approach. The operating principle of this framework is to train the model using a small portion of weakly labeled data (such as point annotations) and a large portion of unlabeled samples. We added our modules to the training process of the PWOOD framework and conducted experiments on the DOTA-v1.0 and DOTA-v1.5 datasets following the training process of PWOOD, setting the proportion of weakly labeled (point labeled) data to 10%, 20%, and 30% respectively.
As shown in Table 7, these results indicate that the approach we adopted in this partially weakly-supervised environment is successful. We conducted ablation experiments and multi-data experiments. In both datasets, our modular approach consistently and significantly outperformed the PWOOD baseline. For example, on the DOTA-v1.0 dataset with only 10% point-labeled data, our method increased the AP50 from 42.35% to 50.67%, an improvement of 8.32% . The performance improvement brought by our method is also robust in cases of higher weak supervision. In the 100% point-labeled (marked as “full”) scenario, our method achieved an AP50 of 64.57%, 8.7% higher than the baseline. It is notable that these improvements were also significant on the more challenging DOTA-v1.5 dataset; for instance, in 10% of the scenarios, it increased from 35.33% to 43.25%, an improvement of 7.92%.
The above experiments have confirmed that our method can be applied in partially weakly-supervised scenarios, significantly improving the performance of the PWOOD framework, whether in the case of only a small amount of weak supervision or with more supervision. This provides a powerful and cost-effective solution for oriented object detection.
5 Conclusion
This paper introduces Point2RBox-v3, an upgraded framework that significantly improves both the quality and utilization efficiency of pseudo labels. We introduce Progressive Label Assignment (PLA), which revitalizes point-supervised FPN and unifies label assignment with fully-supervised paradigms through dynamically generated pseudo labels. To address potential limitations of Voronoi Watershed Loss in object scale learning, we propose Prior-Guided Dynamic Mask Loss (PGDM-Loss), which incorporates Segment Anything Model (SAM) to enhance the quality of pseudo labels.
Experiments yield the following observations: 1) In point-supervised tasks, both the quality and utilization efficiency of pseudo labels are critical factors that can yield substantial performance gains. 2) The label assignment module distributes objects to FPN layers based on their scale-specific receptive fields. Due to the intrinsic spatial approximation of convolutional operations and empirically defined assignment thresholds, this approach exhibits inherent tolerance to pseudo labels imperfections while maintaining detection accuracy. 3) It advances the state of the art by a large amount, achieving 66.09%, 56.86%, 41.28%, 46.40%, 19.60% and 45.96% on the DOTA-v1.0, DOTA-v1.5, DOTA-v2.0, DIOR, STAR and RSAR respectively.